OpenStudy (anonymous):

Please help! Medal awarded. What is the measure of the complementary angle to 1.4 radians? Give an exact answer.

ganeshie8 (ganeshie8):

Complementary angles add up to \(\pi/2\); so you need to solve \(x\) in : \(\large 1.4 + x = \pi/2\)

OpenStudy (anonymous):

ganeshie8, I got decimal and it's asking for radians. i'm a little lost

OpenStudy (ribhu):

@rvc .... hello

rvc (rvc):

hi @ribhu

OpenStudy (anonymous):

can you guys please help

OpenStudy (jhannybean):

Convert 1.4 radians to a fractional representation. \[1.4 ~\text{rad}~ = 1\frac{4}{10}~\pi = \frac{14}{10}\pi = \frac{7\pi}{5}\] And now solve for x.

OpenStudy (jhannybean):

Do you have answer choices?

OpenStudy (anonymous):

thanks so much! I will type them out right now :)